OpenStudy (anonymous):

I have the answer can somone please check me..... If E(t) represents the energy consumed in a househole in watts/ month what are the units of

OpenStudy (anonymous):

\[\int\limits_{a}^{b}E(t)dt\]

OpenStudy (phi):

integration is summing up E* dt (energy * time) if E is in watts/month and time is measured in months , the units are watts/month * month which you can simplify

OpenStudy (anonymous):

I think the answer is watts/month^2

OpenStudy (anonymous):

is that correct

OpenStudy (phi):

\[ \frac{watts}{month} \cdot month \]

OpenStudy (phi):

or , if you like \[ \frac{watts}{month}\cdot \frac{months}{1} \]

OpenStudy (anonymous):

right so then the answer is watts/month^2

OpenStudy (phi):

if it were numbers, for example \[ \frac{4}{2} \cdot \frac{2}{1} \] what do you get ?

OpenStudy (anonymous):

oh are you saying it is just watts is the answer

OpenStudy (phi):

yes, it is just watts \[ \frac{watts}{months} \cdot months = watts \cdot \frac{months}{months} = watts \]

OpenStudy (anonymous):

ok thanks I understnad now months get canceled out thanks